OpenStudy (anonymous):

write an equation of the line: m= -2/3 , contains (-9,4)

OpenStudy (anonymous):

The general equation of a line where we know the slope and one of the points (x1, y1) is: y-(y1) = m(x - x1) So you should get: y - 4 = (-2/3)(x + 9)

OpenStudy (anonymous):

ok thanks